OpenStudy (anonymous):

what is the value of pie(22\7)????

OpenStudy (anonymous):

Do you mean how close is 22/7 to the value of pie? Pie is irrational and goes on forever after the decimal 3.141592654,,,,, 22/7 is close to pi 22/7 = 3.142857143

OpenStudy (anonymous):

\( \large \pi \neq \frac{22}7 \)

OpenStudy (anonymous):

thnx,but isnt it 3.14387???

OpenStudy (anonymous):

No, \( \pi \) is irrational so it's non-terminating and non-recurring , all we can do is approximate to desired digits..

OpenStudy (anonymous):

but specific 22\7

OpenStudy (anonymous):

???

OpenStudy (anonymous):

22\7 is not a good approximation of pi in real mathematics, but works in grade school though.

OpenStudy (anonymous):

Not sure what you are asking 22/7 is not 3.14387.... but it is 3.14287 which is close to pi

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

the history of pi is very interesting on http://www.gap-system.org/~history/HistTopics/Pi_through_the_ages.html